The earth revieved about 1300W/m^2 energy from the sun. Assuming the energy to be in the form of plane polarized monochromatic waves and assuming Normal incidence, calculate




1. The magnitude of the electric filed vector




2. The magnitude of the magnetic field vector

1
Expert's answer
2021-12-27T08:02:50-0500

1. The magnitude of the electric filed vector can be found with the following equation:


"E=\sqrt{2I\sqrt{\frac{\mu_o}{\epsilon_0}}}=990\text{ V/m}."

2. The magnitude of the magnetic filed vector can be found with the following equation:


"H=\sqrt{2I\sqrt{\frac{\epsilon_0}{\mu_o}}}=2.63\text{ A/m}."